Larry was accepted at three different graduate schools, and must choose one. Elite U costs $50,000 per year and did not offer Larry any financial aid. Larry values attending Elite U at $60,000 per year. State College costs $30,000 per year, and offered Larry an annual $10,000 scholarship. Larry values attending State College at $40,000 per year. NoName U costs $20,000 per year, and offered Larry a full $20,000 annual scholarship. Larry values attending NoName at $15,000 per year. Larry maximizes his economic surplus by attending:

a. Elite U.

b. State College.

c. NoName U because he has a full scholarship there.

d. NoName U because the annual cost is only $20,000.
Answer #1

Correct Answer:

B

Working note:

Consumer surplus with Elite U = 60000 - 50000 = $10000

Consumer surplus with State college = 40000 - (30000-10000) = $20000

Consumer surplus with NoName U = 15000 - (20000-20000) = $15000

Here, the surplus is maximum with state college.